OpenStudy (anonymous):

**WILL MEDAL** A crater shaped like a cone has a height of 24 yd and a diameter of 72 yd. What is the volume of the crater? Use 3.14 to approximate pi and express your final answer in hundredths. ___yd3

OpenStudy (jhannybean):

First, what is the volume of a cone?

OpenStudy (anonymous):

Volume is 24?

OpenStudy (jhannybean):

Not quite, the formula for the volume of a cone is : \(\large V=\pi r^2\left(\frac{h}{3}\right)\)

OpenStudy (jhannybean):

you're given the diameter, and dividing the diameter by 2 gives the radius :) \[r=\frac{D}{2}\]

OpenStudy (jhannybean):

\[h=24~yd~,~ \pi= 3.14\] Just plug in these values into your formula to find your volume! :)

OpenStudy (anonymous):

Ok just a sec

OpenStudy (anonymous):

Ok I got 2712.96

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Hmm.. not quite. What did you get for your value of r?

OpenStudy (anonymous):

36

OpenStudy (jhannybean):

Good, me too. What about \(\frac{24}{3} =~?\)

OpenStudy (anonymous):

8

OpenStudy (jhannybean):

Awesome. So now we plug these values into our equation. \[V=(3.14)(36)^2(8)\] And what do you get?

OpenStudy (anonymous):

3255.52

OpenStudy (jhannybean):

I'm getting 32555.52 cm\(^3\) . Are you plugging it in to your calculator properly?

OpenStudy (anonymous):

Oh sorry forgot the cm3 part :/

OpenStudy (anonymous):

and the extra 5

OpenStudy (jhannybean):

Ah, alright :)

OpenStudy (anonymous):

Thank u(: